OpenStudy (dls):

If the tangent to the curve 3xy^2-2x^2y=1 at (1,1) intersect the curve again at (alpha,beta) then find (alpha,beta)

OpenStudy (dls):

\[\Huge 3xy^2-2x^2y=1\]

OpenStudy (primeralph):

find the equation of the line and find the points of intersection.

OpenStudy (dls):

\[\LARGE 3y^2+6xy \frac{dy}{dx}-(4xy+2x^2 \frac{dy}{dx})=0\] \[\LARGE \frac{dy}{dx}=\frac{4xy-3y^2}{6xy-2x^2}\]

OpenStudy (primeralph):

plug in 1,1

OpenStudy (dls):

4-3/6-2 1/4

OpenStudy (primeralph):

Then use y-1 = m(x-1) to find the line equation

OpenStudy (dls):

y-1=1/4(x-1)

OpenStudy (dls):

4y-4=x-1 4y-x-3=0

OpenStudy (primeralph):

Yeah, solve the equations.